OpenStudy (anonymous):

Write the standard equation of a circle with each given radius and center. 1. r = 4; C= (0, 0) 2. r = 1; C = (2, 3) 3. r = 4; C = (-2, 8) Write the standard equation for the circle given the following equations. 4. x2 + y2 - 8x + 7 = 0 5. x2 + y2 + 4x - 6y - 3 = 0

OpenStudy (solomonzelman):

( With (h,K) being the center of the circle AND r being the radius ) \(\color{blue}{ (x-h)^2+(y-k)^2=r^2 }\) plug in your numbers..... \(\color{blue}{ (x-0)^2+(y-0)^2=(4)^2 }\) simplify more.... \(\color{blue}{ x^2+y^2=16 }\) there it is ! can you do the rest based on this example (applying the same formula and Technics ?